Determine the number of molecules of SO2 contained in 49.5 g of SO2.
Answer #1

w= 49.5 g

MB = 64 g = molar mass

NA = avogadro no.= 6.022 x1023

no. of molecules = (w/MB) x NA = (49.5/64) x 6.022 x1023

no. of molecules = 4.658 x 1023